An xray is operated at a pd of 10^5V. Calculate the minimum xray wavelength obtained from the tube.
1
Expert's answer
2020-04-10T08:42:48-0400

"\lambda_{min}=\frac{hc}{eU}=\frac{6.62\cdot 10^{-34}\cdot 3\cdot10^{8}}{1.6\cdot10^{-19}\cdot10^5}=12.4\cdot10^{-12}m"